OpenStudy (anonymous):

square root property (5k+3)^2=12

OpenStudy (anonymous):

k = 1/5 (-3-2 sqrt(3)) k = 1/5 (2 sqrt(3)-3)

OpenStudy (anonymous):

??

OpenStudy (anonymous):

What else do you want to know?

OpenStudy (anonymous):

i just didnt understand at first but i got it